OpenStudy (anonymous):

find the positve value of k for which the given equation as real and equal roots 4x^2-3kx+1=0

OpenStudy (ash2326):

for equal root the determinant D=0 ax^2+bx+c=0 D=b^2-4ac here equation is 4x^2-3kx+1=0 a=4 b=-3k c=1 D=9k^2-4*4*1 D=9k^2-16 D=0 for equal roots 9k^2-16=0 k^2=16/9 k=4/3 or -4/3